A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1890–1899

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

30 November 2021 Version 1.00

Title word cross-reference

12th [Doy96]. 16th [Ano94c]. 17th [Ano94c]. 1894 [Ano94c]. 1895 [Doy96]. 1896 [Ano97d]. 1897 [Ano97g]. 1898 [Ano98h].

20th [Ano97d]. 25th [Doy92].

’96 [Ano96b]. ’97 [Ano97e].

A-Piece [Hal94]. Accessible [Swe94]. Address [Ano94a]. Adhesive [Rei93b, Str94]. Adult [Bum98]. Advancement [Tom97]. Against [Ano94b]. Ain [RW92]. Alaska [Ano98i, Jam94]. Alaskan [Bea90]. Allies [Bea90]. America [Bor93]. American [Ano90b, Ano91d, Ano93c, Ano94c, Ano97d, Ano97g, Ano98h, Doy92, Doy96, Ano90a, Ano91b, Ano93b, Ano94e, Ano96a, Ano97c, Che91, Jam92, Mat90a, Sto98, Swe94]. Among [O’B96].

Angling [Hen93]. Animals [Bir97]. Annual [Ano90b, Ano91d, Ano93c, Ano94c, Ano97d, Ano97h, Ano98h, Ano98k, Doy92, Doy96]. Applied [Ano98d]. Aquacultural [War98]. Aquaria [Sea90]. Aquarium [Ano97d, Doy96]. Aragon [Mun92]. Areas [Ano98j]. Arkansas [Hye90]. Artificial [Hen98, Nev98, Pot90, Sup96]. Artificially [Bum98, Gil92]. Assisted [Ano98e].

Bass [Ano98b, Bow97b]. Battery [Ano97d]. be [GS90]. been [Bum98]. Before [Ano94e]. between [Ano98j]. Biology [Rei97]. Black [Ano98b, Bow97b]. Book [Swe94]. Breeding [Mat91]. Brook [Sup96].

C [Mee97]. Caledonia [Dod96]. Capt. [Ano94a]. Cash [Swe94]. Castle [Ano97d, Doy96]. Certain [Ano98e]. Characteristics [Rei97]. Chinook [Sto94]. chouica [Sto94]. City [Doy92]. Civilizaiton [Ano94b]. Close [Nev98]. Closed [Dic98]. Coast [Ano98i]. Collection [Tit97]. Collins [Ano94a]. Commission [Cul97a, Bea96, Cla91]. Committee [Ano90c, Ano91c, Ano92a, Ano93d, Ano94d, Ano96b, Ano97e, Ano97f, Ano98g]. Community [Ano94e]. Compiled [Swe94]. Concerning [Cul97a, Hen98]. Connection [Ano98f]. Constitution [Ano91a, Ano93a, Ano97a]. Control [Pos92]. Cultivation [Ano92b]. Culture [Ano92b, Ano98d, De 97, GS90, Hen98, Mat96, Pot90, Sea92, Sto94, Tit98]. Culturist [Ano98a].

Danger [Mat90b]. Davis [Gun97]. Decadence [Van96]. Desirability [Sea90]. Desisability [Tit98]. Detroit [Ano97g, Bis90a]. Develop [GS90]. Distributing [Cla91]. Distribution [Bir97, Cla92b, Tom96]. Domesticated [Pag93a]. Done [GS90].

Early [Whi92a]. Eggs [Mat90b, Mat90c, Pag92, Rei93b, Str94, Whi90]. Emery [Gun97]. England [MD93]. Epidemic [O’B96]. Establishment [Mun92, RW92, Sea90]. Estimates [Rei93a]. Europe [Bor93]. Executive [Ano90c, Ano91c, Ano92a, Ano93d, Ano94d, Ano96b, Ano97e, Ano97f, Ano98g]. Exhibit [Hen93, Col93a]. Experiment [Hye90, War98]. Experiments [Pos90, Whi90]. Eyed [Nev97].

Fair [Col91, Col93a, Hen93]. Family [Tom96]. Famine [Ano94b]. Feeding [Cul97b, Sto94]. Fifth [Ano97d]. Fingerlings [Nev92, Whi92b]. First [Doy92]. Fish [Ams96, Ano92b, Ano98a, Ano98d, Ano98f, Ano98i, Bea96, Bis90a, Bow97a, Bum98, Che91, Cla91, Cla92b, Cou92, Dic98, Dod96, Fro97, GS90, Gun92, Hen90, Jam96, Jam97b, Mat90b, Mat96, Pag93a, Pot90, Sto94, Str91, Sup96, Tit98, Tom97]. Fish-Eggs [Mat90b]. Fishculture [Atk94, Bor93, Pos93, Tom93, Whi96b]. Fisheries [Ano90a, Ano90b, Ano91b, Ano91d, Ano93b, Ano93c, Ano94c, Ano94e, Ano96a, Ano97c, Ano97d, Ano97g, Ano98h, Cha92, Che91, Col91, Col93b, Doy92, Doy96, Jam94, Kun93, Mat90a, Pos92, Sto98, Swe94, Whi92a, Col93a, Cul97a]. Fishes [Ano98e, Che92, Hen91, Hun97, Jam92, Jam94, Jam97a, Sea91, Tit98]. Fishing [Gun92]. Fishways [Rog92]. Food [Ano98i, Atk94, Che92, Cou92, Cul97b, Hen91, Hun97, Jam92, Jam94, Jam96, Jam97a, Jam97b, Mat97, Tit98]. Food-Fish [Jam96]. Food-Problem [Atk94]. Ford [Mee97]. Forest [Cul97a]. Fourth [MD93, Doy96]. France [RW92]. Fresh [Rei97, Sto94]. Fry [Cul97b, Mat92, Mat97, Nev92, Whi92b]. Fungus [Dod96]. Further [GS90]. Future [Ano92b].

Game [Ams96, Cul97a, Fro97, Tit98]. Garden [Ano97d, Doy96]. Glance [Ano92b]. Grayling [Bis90b, Hen98]. Great [Sea90, Nev98, Rei93a, Whi92a]. Gremaz [RW92]. Guide [Hen91].

Habits [Hen91, Mat91, Sto94]. Handling [Rei93b, Str94]. Hatched [Bum98]. Hatcheries [Ano98f]. Hatchery [Whi96a]. Hatching [Hye90, Mat94, Pos90, Sup96]. Held [Ano94c, Ano97d, Ano97g, Doy92, Doy96]. Henry [Mee97]. Historical [Ano92b]. History [Whi92a]. Holland [Doy92]. Hon. [Gun97]. Hotel [Ano98h]. House [Ano97g, Doy92].

Identification [Bum98]. Impoverishment [Jam96]. Impregnating [Pag92]. Impregnation [Whi90]. Improved [Mat94]. Industries [Jam96]. Influence [Mat96]. Inland [Ano98j, Bir97]. Inquiry [GS90]. Insurance [Ano94b]. Inter [Jam97a]. Inter-State [Jam97a]. Interests [Jam94].

Jersey [Fro97]. July [Ano98h]. June [Ano97g, Doy96].

Kennerly [Bea91]. Kinds [Ano98e]. Know [Mat93].

Lakes [Ano98j, Bir97, Nev98, Rei93a, Whi92a]. Large [Ano98b]. Large-Mouthed [Ano98b]. Laws [Che91, Jam97b]. Legislate [Kun93]. Legislation [Tit98]. List [Ano96a, Ano97b, Ano97c, Ano98c, Swe94]. Live [Sea91]. Lobster [Mat93]. Local [Ano91c, Ano92a]. Lower [Bir97].

May [Ano94c, Ano97d, Doy92, GS90]. Meeting [Ano90b, Ano91d, Ano93c, Ano94c, Ano97d, Ano97g, Ano98h, Doy92, Doy96, Ano97h, Ano98k]. Members [Ano90a, Ano91b, Ano93b, Ano96a, Ano97b, Ano97c, Ano98c, Swe94]. Method [De 97, Mat94]. Methods [Tit97]. Mich [Ano97g]. Michigan [Ano98f, Bis90b, Pos93]. Microscope [Ano98d]. Millard [Ano98h]. Minutes [Ano93c, Ano94c, Ano97d, Doy92, Doy96]. Modern [Whi96b]. Moral [Whi96b]. Mouth [Bow97b]. Mouthed [Ano98b].

N [Ano97d, Doy96]. National [Sto92]. Nationalism [Cha92]. Natural [Ano98e, Mat97]. Neb [Ano98h]. Nebraska [O’B96]. Need [Cha94]. Needed [Pag93b]. Nineteenth [Ano90b]. Non [Sto94]. Non-Feeding [Sto94]. North [Bor93]. Northville [Cla91]. Norwegian [Col93b]. Notes [Ano98f].

Observations [Hen98, Whi96b]. Occurrence [Kun93]. Officers [Ano90c, Ano91c, Ano92a, Ano93d, Ano94d, Ano96b, Ano97e, Ano97f, Ano98g]. Ohio [Gun92]. Omaha [Ano98h]. Oncorhynchus [Sto94]. Organization [Tit98]. Origin [Pot90, Sto98]. other [Jam92]. Our [Ano94b, Van96].

Pacific [Ano98i]. Paper [Ano94e]. Park [Ano97d, Sto92]. past [Ano92b, GS90]. Pearls [Kun93]. Penny [Hal94]. Perch [Mat90c, Mat91]. Phases [Whi96b]. Philadelphia [Ano94c]. Piece [Hal94]. Piedra [Mun92]. Pike [Mat90c, Nev97]. Pike-Perch [Mat90c]. Pikeperch [Whi90]. Piscicultural [Mun92, RW92]. Plankton [Rei93a]. Plant [Pag93b]. Planting [Mat92, Nev92, Sup96]. Plants [Bir97]. Pond [De 97]. Potter [Gun97]. Practically [Ano98d]. Preliminary [Hen98]. Present [GS90, Sea92]. Preserve [Kun93]. Pressing [Cha94]. Problem [Atk94, Cul97b]. Proceedings [Mat90a, Ano97g, Ano98h]. Procurement [Tit98]. Production [Bow97a, Tom97]. Promote [GS90]. Promotion [Tit98]. Propagation [Bow97b, Jam97b, Nev98, Tit98]. Protection [Ams96, Ano98i, Bow97a, Dic98, Fro97, Hen90, Jam97a, Jam97b, Sto94, Tit98]. Protective [Che91]. Public [Sea90]. Published [Swe94].

quinnat. [Sto94].

Railroads [Mat96]. Rainbow [Pag92]. Read [Ano94e]. Reared [Gil92]. Rearing [Cla91, Cla92b, Hye90]. Regular [Ano97g]. Relation [Ano98j, Che91]. Relations [Ano94e, Sup96]. Report [Mun92]. Reproducing [Cou92]. Reproduction [Ano98e]. Results [Sup96]. Review [Bor93]. Rolls [Swe94]. Russell [Ano97g].

S [Cla91]. S. [Mat90c]. Saibling [MD93]. Sale [Gil92, Pag93a]. Salmo [Sto94]. Salmon [Bea90, Bea91, Jam92, Sto92, Sto94]. Salvelinus [MD93]. Science [GS90]. Season [Dic98, Nev98]. Second [Ano93c]. Self [Cou92]. Self-Reproducing [Cou92]. Seventh [Ano98h, Ano98k]. Shad [Hal94]. Shall [Kun93]. Siskiwit [Swe90]. Sixth [Ano97h]. Small [Bow97b]. Smelts [Mat94]. Society [Ano94c, Ano97d, Ano97g, Ano98h, Doy92, Doy96, Ano90a, Ano90b, Ano91b, Ano91d, Ano93b, Ano93c, Ano94e, Ano96a, Ano97c, Che91, Mat90a, Sto98, Swe94]. Solved [Cul97b]. Some [Hen98, Pos90, Rei93a, Rei97, Whi96b]. Spain [Mun92]. Spawn [Tit97]. Specialist [Tom93]. State [Cha92, Jam97a, Jam97b, Pos92, Tit98]. States [Bea96, Ano98f, Kun93, Pot90, Sea90]. Station [Cla91, Bis90a]. Stations [War98]. Statistical [Bor93]. Status [Sea92]. Streams [Van96]. Sturgeon [Pos90]. Sufficient [Ano94b]. Suggestion [Tom93]. Sunapee [MD93]. Supply [Ano98i]. Susquehanna [Cla92a].

Teeth [Hen91]. Temperature [Ano98j]. their [Bea90, Hen91, War98]. Third [Ano94c]. Thursday [Ano94c]. Transit [Mat90b]. Transplanting [Str91]. Transportation [Sea91]. Trout [Cla91, Cul97b, Gil92, Hye90, Mat92, Mat97, O’B96, Pag92, Sea92, Sup96, Tit97, Tom96, Van96]. Twentieth [Ano91d]. Twenty [Ano98h, Ano93c, Ano94c, Ano97d, Ano97h, Ano98k, Doy92, Doy96]. Twenty-Fifth [Ano97d]. Twenty-First [Doy92]. Twenty-Fourth [Doy96]. Twenty-Second [Ano93c]. Twenty-Seventh [Ano98h, Ano98k]. Twenty-Sixth [Ano97h]. Twenty-Third [Ano94c].

U [Cla91]. Uniform [Jam97b]. United [Ano98f, Bea96, Kun93, Pot90, Sea90]. Utility [Tit97].

Variety [MD93]. versus [Ano98e, Nev98]. Vertical [Bir97]. Vitreum [Mat90c]. vs [Nev92, Whi92b].

Wall [Nev97]. Wall-Eyed [Nev97]. Waste [Hun97]. Water [Ano98j]. Waters [Ano92b, Rei97, Sto94]. Wednesday [Ano94c, Ano97d, Doy92, Doy96]. were [Hal94]. Where [Pag93b]. White [Bis90a, Sup96]. White-Fish [Bis90a, Sup96]. Wild [Tit97]. Work [Bea96, Cul97a, Rei97, War98]. World [Col91, Col93a, Hen93].

Yearlings [Mat92, Pag93b]. York [Doy92, Doy96, Cul97a]. Young [Cou92].

References

Amsden:1896:GFP

[Ams96] Frank J. Amsden. Game and fish protection. Transactions of the American Fisheries Society, 24(1):100–108, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1890:MAF

[Ano90a] Anonymous. Members of the American Fisheries Society. Transactions of the American Fisheries Society, 19(1):116–122, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1890:NAM

[Ano90b] Anonymous. Nineteenth Annual Meeting of the American Fisheries Society. Transactions of the American Fisheries Society, 19(1):3–15, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1890:OEC

[Ano90c] Anonymous. Officers for 1890–91 and Executive Committee. Transactions of the American Fisheries Society, 19(1):2, January 1890. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1891:C

[Ano91a] Anonymous. Constitution. Transactions of the American Fisheries Society, 20(1):61–62, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1891:MAF

[Ano91b] Anonymous. Members of the American Fisheries Society. Transactions of the American Fisheries Society, 20(1):63–69, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1891:OEC

[Ano91c] Anonymous. Officers for 1891–92, Executive Committee and Local Committee. Transactions of the American Fisheries Society, 20(1):2, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1891:TAM

[Ano91d] Anonymous. Twentieth Annual Meeting of the American Fisheries Society. Transactions of the American Fisheries Society, 20(1):3–24, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1892:OEC

[Ano92a] Anonymous. Officers for 1892–93, Executive Committee and Local Committee. Transactions of the American Fisheries Society, 21(1):2, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1892:PFF

[Ano92b] Anonymous. The past and future of fish culture: an historical glance at the cultivation of the waters. Transactions of the American Fisheries Society, 21(1):126–127, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1893:C

[Ano93a] Anonymous. Constitution. Transactions of the American Fisheries Society, 22(1):167–168, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1893:MAF

[Ano93b] Anonymous. Members of the American Fisheries Society. Transactions of the American Fisheries Society, 22(1):169–178, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1893:MTS

[Ano93c] Anonymous. Minutes of the Twenty-Second Annual Meeting of the American Fisheries Society. Transactions of the American Fisheries Society, 22(1):3–14, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1893:OEC

[Ano93d] Anonymous. Officers for 1893–94 and Executive Committee. Transactions of the American Fisheries Society, 22(1):2, January 1893. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1894:ACC

[Ano94a] Anonymous. Address of Capt. Collins. Transactions of the American Fisheries Society, 23(1):84–92, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1894:OCS

[Ano94b] Anonymous. Has our civilizaiton sufficient insurance against famine? Transactions of the American Fisheries Society, 23(1):93–109, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1894:MTT

[Ano94c] Anonymous. Minutes of the Twenty-Third Annual Meeting of the American Fisheries Society: Held in Philadelphia, Wednesday, May 16th and Thursday, May 17th, 1894. Transactions of the American Fisheries Society, 23(1):3–17, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1894:OEC

[Ano94d] Anonymous. Officers for 1894–95 and Executive Committee. Transactions of the American Fisheries Society, 23(1):2, January 1894. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1894:RCF

[Ano94e] Anonymous. Relations of the community to the fisheries: (a paper read before the American Fisheries Society, 1894.). Transactions of the American Fisheries Society, 23(1):35–37, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1896:LMA

[Ano96a] Anonymous. List of members of the American Fisheries Society. Transactions of the American Fisheries Society, 24(1):115–123, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1896:OEC

[Ano96b] Anonymous. Officers for 1895–’96 and Executive Committee. Transactions of the American Fisheries Society, 24(1):2, January 1896. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1897:C

[Ano97a] Anonymous. Constitution. Transactions of the American Fisheries Society, 26(1):131–132, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1897:LM

[Ano97b] Anonymous. List of members. Transactions of the American Fisheries Society, 26(1):132–136, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1897:LMA

[Ano97c] Anonymous. List of members of the American Fisheries Society. Transactions of the American Fisheries Society, 25(1):146–153, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1897:MTF

[Ano97d] Anonymous. Minutes of the Twenty-Fifth Annual Meeting of the American Fisheries Society: Held in Battery Park Aquarium, Castle Garden, N. Y., On Wednesday, May 20th, 1896. Transactions of the American Fisheries Society, 25(1):5–36, January 1897. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1897:OECa

[Ano97e] Anonymous. Officers for 1896–’97 and Executive Committee. Transactions of the American Fisheries Society, 25(1):3, January 1897. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1897:OECb

[Ano97f] Anonymous. Officers for 1897–98 and Executive Committee. Transactions of the American Fisheries Society, 26(1):2, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1897:PAF

[Ano97g] Anonymous. Proceedings of American Fisheries Society at its Regular Meeting Held at the Russell House, Detroit, Mich., June 17, 18 and 19, 1897. Transactions of the American Fisheries Society, 26(1):5– 24, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1897:TSA

[Ano97h] Anonymous. Twenty-Sixth Annual Meeting. Transactions of the American Fisheries Society, 26(1):50–57, July 1897. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:FC

[Ano98a] Anonymous. The fish culturist. Transactions of the American Fisheries Society, 27(1):93–99, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:LMB

[Ano98b] Anonymous. Large-mouthed black bass. Transactions of the American Fisheries Society, 27(1):83–88, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:LM

[Ano98c] Anonymous. List of members. Transactions of the American Fisheries Society, 27(1):136–140, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:MPA

[Ano98d] Anonymous. The microscope as practically applied to fish culture. Transactions of the American Fisheries Society, 27(1):88–93, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:NVA

[Ano98e] Anonymous. Natural versus assisted reproduction of certain kinds of fishes. Transactions of the American Fisheries Society, 27(1):46–56, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:NCU

[Ano98f] Anonymous. Notes in connection with the United States fish hatcheries in Michigan. Transactions of the American Fisheries Society, 27(1):27–32, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:OEC

[Ano98g] Anonymous. Officers for 1898–99 and Executive Committee. Transactions of the American Fisheries Society, 27(1):2, January 1898. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:PTS

[Ano98h] Anonymous. Proceedings of Twenty-Seventh Annual Meeting of American Fisheries Society at the Hotel Millard, Omaha, Neb., July 20, 21 and 22, 1898. Transactions of the American Fisheries Society, 27(1):5–17, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:PFF

[Ano98i] Anonymous. Protection of the food fish supply on the Pacific Coast and in Alaska. Transactions of the American Fisheries Society, 27(1): 112–120, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:RBA

[Ano98j] Anonymous. The relation between the areas of inland lakes and the temperature of the water. Transactions of the American Fisheries Society, 27(1):99–105, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1898:TSA

[Ano98k] Anonymous. Twenty-Seventh Annual Meeting. Transactions of the American Fisheries Society, 27(1):61–70, January 1898. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1894:FPF

[Atk94] Chas G. Atkins. The food-problem in fishculture. Transactions of the American Fisheries Society, 23(1):58–66, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1890:AST

[Bea90] Tarleton H. Bean. The Alaskan salmon and their Allies. Transactions of the American Fisheries Society, 19(1):49–66, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1891:KS

[Bea91] Tarleton H. Bean. Kennerly’s salmon. Transactions of the American Fisheries Society, 20(1):41–46, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1896:WUS

[Bea96] Tarleton H. Bean. The work of the United States Fish Commission. Transactions of the American Fisheries Society, 24(1):75–82, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Birge:1897:VDL

[Bir97] E. A. Birge. Vertical distribution of the lower plants and animals in the inland lakes. Transactions of the American Fisheries Society, 26(1):25– 40, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bissell:1890:DWF

[Bis90a] J. H. Bissell. The Detroit White-Fish Station. Transactions of the American Fisheries Society, 19(1):16–20, January 1890. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bissell:1890:GM

[Bis90b] John H. Bissell. Grayling in Michigan. Transactions of the American Fisheries Society, 19(1):27–29, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Borodine:1893:SRF

[Bor93] N. Borodine. Statistical review of fishculture in Europe and North America. Transactions of the American Fisheries Society, 22(1):104– 112, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:1897:FPF

[Bow97a] Seymour Bower. Fish protection and fish production. Transactions of the American Fisheries Society, 26(1):58–65, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:1897:PSM

[Bow97b] Seymour Bower. The propagation of small mouth black bass. Transactions of the American Fisheries Society, 25(1):127–136, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bumpus:1898:IAF

[Bum98] H. C. Bumpus. The identification of adult fish that have been artificially hatched. Transactions of the American Fisheries Society, 27 (1):70–83, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chamberlayne:1892:NSF

[Cha92] C. A. Chamberlayne. Nationalism in state fisheries. Transactions of the American Fisheries Society, 21(1):187–196, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chamberlayne:1894:PN

[Cha94] Charles F. Chamberlayne. The pressing need. Transactions of the American Fisheries Society, 23(1):110–118, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cheney:1891:RAF

[Che91] A. N. Cheney. Relation of the American Fisheries Society to protective fish laws. Transactions of the American Fisheries Society, 20(1):47–51, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cheney:1892:FF

[Che92] A. Nelson Cheney. Food for fishes. Transactions of the American Fisheries Society, 21(1):22–32, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1891:RDT

[Cla91] Frank N. Clark. Rearing and distributing trout at the Northville Station, U. S. Fish Commission. Transactions of the American Fisheries Society, 20(1):30–33, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clapp:1892:S

[Cla92a] A. F. Clapp. The Susquehanna. Transactions of the American Fisheries Society, 21(1):135–138, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1892:RFD

[Cla92b] Frank N. Clark. Rearing fish for distribution. Transactions of the American Fisheries Society, 21(1):78–81, January 1892. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:1891:FWF

[Col91] J. W. Collins. The fisheries at the World’s Fair. Transactions of the American Fisheries Society, 20(1):33–41, January 1891. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:1893:FEW

[Col93a] Capt J. W. Collins. The Fisheries Exhibit at the World’s Fair. Transactions of the American Fisheries Society, 22(1):122–129, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:1893:NF

[Col93b] Capt J. W. Collins. The Norwegian fisheries. Transactions of the American Fisheries Society, 22(1):60–70, January 1893. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Counsul:1892:SRF

[Cou92] Frank H. Mason Counsul. Self-reproducing food for young fish. Transactions of the American Fisheries Society, 21(1):58–77, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culturist:1897:CWF

[Cul97a] A. N. Cheney State Fish Culturist. Concerning the work of the Fisheries, Game, and Forest Commission of the State of New York. Transactions of the American Fisheries Society, 25(1):112–120, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culturist:1897:FTF

[Cul97b] S. E. Land Wyoming Fish Culturist. Feeding trout fry, or the food problem solved. Transactions of the American Fisheries Society, 26 (1):128–130, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeBellesme:1897:NMP

[De 97] Jousset De Bellesme. New method of pond culture. Transactions of the American Fisheries Society, 25(1):69–94, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dickenson:1898:PFC

[Dic98] F. B. Dickenson. The protection of fish and a closed season. Transactions of the American Fisheries Society, 27(1):32–46, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dodge:1896:FFC

[Dod96] Charles Wright Dodge. Fish fungus at Caledonia. Transactions of the American Fisheries Society, 24(1):109–114, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doyle:1892:MTF

[Doy92] Edward P. Doyle. Minutes of the Twenty-First Annual Meeting of the American Fisheries Society: Held in the City of New York. Wednesday, May 25th, 1892, at the Holland House. Transactions of the American Fisheries Society, 21(1):3–14, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doyle:1896:MTF

[Doy96] Edward P. Doyle. Minutes of the Twenty-Fourth Annual Meeting of the American Fisheries Society: Held in New York Aquarium, Castle Garden, N. Y., on Wednesday, June 12th, 1895. Transactions of the American Fisheries Society, 24(1):5–16, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frothingham:1897:FGP

[Fro97] H. P. Frothingham. Fish and game protection in New Jersey. Transactions of the American Fisheries Society, 25(1):137–145, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gilbert:1892:SAR

[Gil92] W. L. Gilbert. Sale of artificially reared trout. Transactions of the American Fisheries Society, 21(1):115–118, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gay:1890:PPF

[GS90] John Gay and Wm P. Seal. The past and present of fish culture, with an inquiry as to what may be done to further promote and develop the science. Transactions of the American Fisheries Society, 19(1):66–79, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunckel:1892:FFO

[Gun92] John E. Gunckel. Fish and fishing in Ohio. Transactions of the American Fisheries Society, 21(1):15–21, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunckel:1897:HED

[Gun97] J. K. Gunckel. Hon. Emery Davis Potter. Transactions of the American Fisheries Society, 25(1):37–47, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hallock:1894:WSW

[Hal94] Charles Hallock. When shad were a penny a-piece. Transactions of the American Fisheries Society, 23(1):18–21, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1890:FP

[Hen90] James A. Henshall. Fish protection. Transactions of the American Fisheries Society, 19(1):79–84, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1891:TFG

[Hen91] James A. Henshall. On the teeth of fishes as a guide to their food habits. Transactions of the American Fisheries Society, 20(1):24–30, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1893:AEW

[Hen93] James A. Henshall. The angling exhibit at the World’s Fair. Transactions of the American Fisheries Society, 22(1):129–131, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1898:SPO

[Hen98] J. A. Henshall. Some preliminary observations concerning the artificial culture of the grayling. Transactions of the American Fisheries Society, 27(1):105–111, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntington:1897:WFF

[Hun97] L. D. Huntington. Waste of food fishes. Transactions of the American Fisheries Society, 25(1):121–126, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hyerson:1890:ETH

[Hye90] P. P. B. Hyerson. Experiment in trout hatching and rearing in Arkansas. Transactions of the American Fisheries Society, 19(1):46–48, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1892:ASO

[Jam92] Bushrod W. James. American salmon and other food fishes. Transactions of the American Fisheries Society, 21(1):139–148, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1894:AFF

[Jam94] Bushrod W. James. Alaska’s food fishes and the interests of its fisheries. Transactions of the American Fisheries Society, 23(1):67–76, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1896:IFF

[Jam96] Bushrod W. James. Impoverishment of the food-fish industries. Transactions of the American Fisheries Society, 24(1):36–48, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1897:ISP

[Jam97a] Bushrod W. James. Inter-state protection of food fishes. Transactions of the American Fisheries Society, 25(1):95–111, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1897:SLU

[Jam97b] Bushrod W. James. State laws for the uniform protection and propagation of food fish. Transactions of the American Fisheries Society, 26(1):113–125, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kunz:1893:OPU

[Kun93] George Frederic Kunz. On the occurrence of pearls in the United States, and shall we legislate to preserve the fisheries. Transactions of the American Fisheries Society, 22(1):16–34, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1890:AFS

[Mat90a] Fred Mather. The American Fisheries Society and its proceedings. Transactions of the American Fisheries Society, 19(1):87–115, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1890:DFE

[Mat90b] Fred Mather. Danger to fish-eggs in transit. Transactions of the American Fisheries Society, 19(1):43–45, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1890:EPP

[Mat90c] Fred Mather. Eggs of pike-perch — S. Vitreum. Transactions of the American Fisheries Society, 19(1):15–16, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1891:BHP

[Mat91] Fred Mather. Breeding habits of the perch. Transactions of the American Fisheries Society, 20(1):51–55, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1892:PTF

[Mat92] Fred Mather. Planting trout as fry or yearlings. Transactions of the American Fisheries Society, 21(1):86–93, January 1892. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1893:WWK

[Mat93] Fred Mather. What we know of the lobster. Transactions of the American Fisheries Society, 22(1):161–166, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1894:IMH

[Mat94] Fred Mather. Improved method of hatching smelts. Transactions of the American Fisheries Society, 23(1):77–83, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1896:IRF

[Mat96] Fred Mather. The influence of railroads on fish culture. Transactions of the American Fisheries Society, 24(1):17–27, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1897:NFT

[Mat97] Fred Mather. Natural food for trout fry. Transactions of the American Fisheries Society, 25(1):48–68, January 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

M:1893:SSF

[MD93] John D. Quackenbos A. M. and M. D. The Sunapee saibling: a fourth New England variety of Salvelinus. Transactions of the American Fisheries Society, 22(1):34–51, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1897:HCF

[Mee97] William E. Meehan. Henry C. Ford. Transactions of the American Fisheries Society, 26(1):66–72, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muntadas:1892:RPE

[Mun92] F. Muntadas. Report on the piscicultural establishment of Piedra, Aragon, Spain. Transactions of the American Fisheries Society, 21 (1):52–57, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nevin:1892:PFV

[Nev92] James Nevin. Planting fry vs. planting fingerlings. Transactions of the American Fisheries Society, 21(1):81–86, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nevin:1897:WEP

[Nev97] James Nevin. Wall-eyed pike. Transactions of the American Fisheries Society, 26(1):126–127, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nevin:1898:APV

[Nev98] Mr. James Nevin. Artificial propagation versus a close season for the Great Lakes. Transactions of the American Fisheries Society, 27(1): 17–27, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OBrien:1896:EAT

[O’B96] M. E. O’Brien. Epidemic among trout in Nebraska. Transactions of the American Fisheries Society, 24(1):52–58, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:1892:IER

[Pag92] Wm F. Page. Impregnating eggs of the rainbow trout. Transactions of the American Fisheries Society, 21(1):179–186, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:1893:SDF

[Pag93a] W. F. Page. Sale of domesticated fish. Transactions of the American Fisheries Society, 22(1):159–161, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:1893:PYW

[Pag93b] W. M. F. Page. Plant yearlings where needed. Transactions of the American Fisheries Society, 22(1):71–103, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Post:1890:SSE

[Pos90] Hoyt Post. The sturgeon; some experiments in hatching. Transactions of the American Fisheries Society, 19(1):36–40, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Post:1892:SCS

[Pos92] Hovt Post. State control of state fisheries. Transactions of the American Fisheries Society, 21(1):118–125, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Post:1893:FM

[Pos93] Hoyt Post. Fishculture in Michigan. Transactions of the American Fisheries Society, 22(1):132–154, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Potter:1890:OAF

[Pot90] E. D. Potter. The origin of artificial fish culture in the United States. Transactions of the American Fisheries Society, 19(1):41–43, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reighard:1893:SPE

[Rei93a] J. E. Reighard. Some plankton estimates in the Great Lakes. Transactions of the American Fisheries Society, 22(1):112–122, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reighard:1893:HAE

[Rei93b] Jacob Reighard. On the handling of adhesive eggs. Transactions of the American Fisheries Society, 22(1):51–60, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reighard:1897:SCR

[Rei97] Jacob Reighard. Some characteristics of recent work on the biology of fresh waters. Transactions of the American Fisheries Society, 26(1):41– 50, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rogers:1892:F

[Rog92] W. H. Rogers. Fishways. Transactions of the American Fisheries Society, 21(1):127–135, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raveret-Wattell:1892:PEG

[RW92] C. Raveret-Wattell. The piscicultural establishment at Gremaz (Ain), France. Transactions of the American Fisheries Society, 21(1):45–51, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seal:1890:DEG

[Sea90] William P. Seal. On the desirability of the establishment, of great public aquaria in the United States. Transactions of the American Fisheries Society, 19(1):20–26, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seal:1891:TLF

[Sea91] William P. Seal. Transportation of live fishes. Transactions of the American Fisheries Society, 20(1):55–60, January 1891. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seal:1892:PST

[Sea92] Wm P. Seal. The present status of trout culture. Transactions of the American Fisheries Society, 21(1):33–45, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1892:NSP

[Sto92] Livingston Stone. A national salmon park. Transactions of the American Fisheries Society, 21(1):149–162, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1894:FCF

[Sto94] Livingston Stone. Fish culture and fish protection: The Chinook salmon. (Oncorhynchus chouica; Salmo quinnat.) its non-feeding habits in fresh waters. Transactions of the American Fisheries Society, 23(1): 26–34, January 1894. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1898:OAF

[Sto98] Livingston Stone. The origin of the American Fisheries Society. Transactions of the American Fisheries Society, 27(1):56–61, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranahan:1891:TF

[Str91] J. J. Stranahan. Transplanting fish. Transactions of the American Fisheries Society, 20(1):46–47, January 1891. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stranahan:1894:HAE

[Str94] J. J. Stranahan. The handling of adhesive eggs. Transactions of the American Fisheries Society, 23(1):22–25, January 1894. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Supt:1896:AHW

[Sup96] Seymour Bower Supt. The artificial hatching of white-fish and brook trout, and the relations of planting to results. Transactions of the American Fisheries Society, 24(1):89–99, January 1896. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sweeny:1890:S

[Swe90] R. O. Sweeny, Sr. The Siskiwit. Transactions of the American Fisheries Society, 19(1):84–86, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sweeny:1894:LMA

[Swe94] R. O. Sweeny, Sr. List of members of the American Fisheries Society: Compiled from cash book and accessible published rolls. Transactions of the American Fisheries Society, 23(1):119–128, January 1894. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1897:WTS

[Tit97] J. W. Titcomb. Wild trout spawn; methods of collection and utility. Transactions of the American Fisheries Society, 26(1):73–92, July 1897. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1898:DSO

[Tit98] Hon. John W. Titcomb. Desisability of state organization for the promotion of fish culture and for the procurement of state legislation for the propagation and protection of food and game fishes. Transactions of the American Fisheries Society, 27(1):120–125, January 1898. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tomlin:1893:SSF

[Tom93] W. David Tomlin. A suggestion: The specialist in fishculture. Transactions of the American Fisheries Society, 22(1):154–159, January 1893. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tomlin:1896:DTF

[Tom96] W. D. Tomlin. The distribution of the trout family. Transactions of the American Fisheries Society, 24(1):49–51, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tomlin:1897:AFP

[Tom97] W. D. Tomlin. Advancement in fish production. Transactions of the American Fisheries Society, 26(1):93–112, July 1897. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanCleef:1896:DOT

[Van96] J. S. Van Cleef. Decadence of our trout streams. Transactions of the American Fisheries Society, 24(1):28–35, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1898:AES

[War98] Henry B. Ward. Aquacultural experiment stations and their work. Transactions of the American Fisheries Society, 27(1):125–135, January 1898. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitaker:1890:EIP

[Whi90] Herschel Whitaker. Experiments in the impregnation of pikeperch eggs. Transactions of the American Fisheries Society, 19(1):30–36, January 1890. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitaker:1892:EHF

[Whi92a] Herschel Whitaker. Early history of the fisheries on the Great Lakes. Transactions of the American Fisheries Society, 21(1):163–179, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitaker:1892:FVF

[Whi92b] Herschel Whitaker. Fry vs. fingerlings. Transactions of the American Fisheries Society, 21(1):94–114, January 1892. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitaker:1896:NH

[Whi96a] Herschel Whitaker. A new hatchery. Transactions of the American Fisheries Society, 24(1):83–88, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitaker:1896:SOM

[Whi96b] Herschel Whitaker. Some observations on the moral phases of modern fishculture. Transactions of the American Fisheries Society, 24(1):59– 74, January 1896. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).